three horses calle A,B,C are all racing aginst each other the probability that A will win is 1/7 the probability that B will win is 1/3. What is the probability that c will win?? Also what is the odds that C will win 11 to___?
Answer by stanbon(75887)

Three horses called A,B,C are all racing aginst each other.
The probability that A will win is 1/7 the probability that B will win is 1/3. What is the probability that C will win??
---
P(C wins) = 1-[(1/3)+(1/7)] = 1-[10/21] = 11/21
---------------------------------
Also what is the odds that C will win 11 to 10?
Odds C wind = P(win)/P(lose) = (11/21)/(10/21) = 11:10
========================
Cheers,
Stan H.
========================